[Hallc_running] [New Logentry] RC Daily Update

bduran at jlab.org bduran at jlab.org
Mon Feb 17 10:30:02 EST 2020

Logentry Text:
On Saturday at 6:30 pm, we took an opportunistic access while the accelerator was down due to a pass change for Hall B. During the access, ¼ wave plates were adjusted using the absorption spectrum to attempt to fix the low target polarization issue. Around 10:20 pm, MCC finished the pass change and were ready to deliver the beam. We’re done with the laser work and ready for the hall close up around 11:10 pm. While we’re doing the beam checkout, the target group did some NMR measurements and unfortunately didn’t see an increase in the polarization which made them to decide having another access. All the work performed by the target group is explained in details by Junhao in this entry (https://logbooks.jlab.org/entry/3784881). 

We came back to beam permit around 2:30 am on Sunday and proceeded with the beam check out procedure. The first harp scans (https://logbooks.jlab.org/entry/3784870 https://logbooks.jlab.org/entry/3784891) gave wide beam width, so we let MCC to tune the beam. We had the results of second harp scans (< 0.5 mm) at 3:00 am (https://logbooks.jlab.org/entry/3784899) and asked for beam to take carbon optics data as the target polarization was low for the production running. At 7:00 am HMS Dipole threw an interlock, but the reset was successful (https://logbooks.jlab.org/entry/3785004).
Two polarization measurements, at 5:20 am and 8:30 am on Sunday morning gave 33.6% and 36.3% longitudinal polarization, respectively. At 8:30 am we switched to transverse spin direction and NMR done at 10:42 am indicated 40% transverse polarization (https://logbooks.jlab.org/entry/3785111). We have been taking production data since then.
The most recent NMR measurements: 43.72% longitudinal (https://logbooks.jlab.org/entry/3785628) and 49.63% transverse (https://logbooks.jlab.org/entry/3785781)

 Low SHMS LN2 levels (~39%) were noticed by the Sunday swing shiftt crew (https://logbooks.jlab.org/entry/3785541) and it seems it’s been fluctuating   since February 14th (https://logbooks.jlab.org/entry/3785570).

TO of the Sunday owl shift reported an issue with the field correction coil readback which has been causing alarms. (https://logbooks.jlab.org/entry/3785664)

We will have an RC meeting today at 4:15 pm at the second floor of the counting house. Here is how to connect remotely:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial:
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3785950
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200217/2a503415/attachment.html>

More information about the Hallc_running mailing list